A certain airplane has two independent alternators to provide electrical power. The probability that a given alternator will fail on a 1-hour flight is .02. What is that (a) both will fail (b) Neither will fail? (c) One or the other will fail? Show all steps carefully?
Found 2 solutions by stanbon, nickstud:
Answer by stanbon(48568)

A certain airplane has two independent alternators to provide electrical power. The probability that a given alternator will fail on a 1-hour flight is .02. What is that
(a) P(both will fail) = 0.02^2 0.0004
-----------
(b) P(Neither will fail) = 0.08^2 = 0.0064
-----------
(c) P(One or the other will fail) = 1-(0.0004+0.0064) = 0.9932
====================
Cheers,
Stan H.

Answer by nickstud(1)

P(A)=.02 P(B)=.02 A=Alternator #1 B=Alternator #2
a)P(A ∩ B) = P(A)P(B) = (.02)(.02) = .0004
b)(.98)(.98) = 0.9604
c)[P(alt 1 fails) * P(alt 2 doesn't fail)] + [P(alt 1 doesn't fail) * P(alt 2 fails)] = [ .02 * .98] + [.98 * .02] = .0392